OpenStudy (liizzyliizz):

limit of |x-2\/x-2 as x approaches 2. would the limit be 0?

myininaya (myininaya):

so it looks like your function is a little messed up

OpenStudy (anonymous):

not clear to me what the function is.

OpenStudy (liizzyliizz):

ill write it out better hold on

myininaya (myininaya):

\[\lim_{x \rightarrow 2}\frac{|x-2|}{x-2}\] ?

OpenStudy (anonymous):

aaah i thought as much. no limit

OpenStudy (liizzyliizz):

yeah you are correcy myinninaya.

OpenStudy (liizzyliizz):

correct*

OpenStudy (anonymous):

this is a piecewise function. it is 1 if x > 2, -1 if x < 2

myininaya (myininaya):

\[x>2=>|x-2|=x-2, x<2=>|x-2|=-(x-2)\]

OpenStudy (anonymous):

and so the limit from the right is 1, the limit from the left is -1, and since those numbers are not the same the limit does not exist

myininaya (myininaya):

\[\lim_{x \rightarrow 2^-}\frac{-(x-2)}{x-2}=-1\]

myininaya (myininaya):

\[\lim_{x \rightarrow 2^+}\frac{x-2}{x-2}=1\]

OpenStudy (liizzyliizz):

oh ok yeah i made a table and i got those results, but i wasnt sure if there was a limit or not. i have to write why, but im not quite sure of the reason.

myininaya (myininaya):

\[\lim_{x \rightarrow 2}\frac{|x-2|}{x-2} DNE\]

OpenStudy (anonymous):

\[f(x) = \frac{|x-2|}{x-2} = \left\{\begin{array}{rcc} 1 & \text{if} & x \geq 2 \\ -1& \text{if} & x < 2 \end{array} \right.\]

OpenStudy (anonymous):

tada!

myininaya (myininaya):

do you know if we have |x|, then: x>0 => |x|=x x<0 => |x|=-x x=0 => |x|=0

myininaya (myininaya):

why is it writing everything on one line :(

OpenStudy (liizzyliizz):

hmmm. ?